OpenStudy (anonymous):

What is the slope of the line passing through (1, 2) and (3, 8)?

OpenStudy (shubhamsrg):

what if i say 4 is the ans..?? would you trust me ?

OpenStudy (shubhamsrg):

:D

OpenStudy (anonymous):

mmmm maybe you had a mustach and some sunglasses

OpenStudy (shubhamsrg):

whats the problem even if i had it ?

OpenStudy (anonymous):

so the slope is 3 jackelynn

OpenStudy (anonymous):

you would look more appealing to trust

OpenStudy (shubhamsrg):

and i thought only i was weird!! lol..

OpenStudy (saranya):

The answer is simple, (8-3)/(2-(1)) = 5/1 or 5. i guess :D

OpenStudy (shubhamsrg):

well theres the female version of me! lol..

OpenStudy (anonymous):

maybe you eat alot of ham

OpenStudy (shubhamsrg):

hmm,,you were trying to make some taunt or something like that isnt it? o.O

OpenStudy (jackellyn):

@saranya look at the points again. 8 and 2 are y and 3 and 1 are x.

OpenStudy (shubhamsrg):

she was just being sarcastic @jackellyn (i hope :P )

OpenStudy (saranya):

oh yeah juz realized t sry for the mistake ppl :( then it must be 8-2/ 3-1 =3

OpenStudy (anonymous):

hamm sometimes i dont understand what youre sayin are mexican?

OpenStudy (anonymous):

are you mexican? hamm?

OpenStudy (saranya):

Shubham yu r too crazy :P

OpenStudy (shubhamsrg):

try oxford dictionary buddy..

OpenStudy (saranya):

seriously :P

OpenStudy (shubhamsrg):

lol..just in some weird mood! :P

OpenStudy (saranya):

lol :P GO n find some good job lol :P

OpenStudy (anonymous):

yeahh go sell hotdogs in nyc im tellin you

OpenStudy (shubhamsrg):

and you're gonna finance me?

OpenStudy (anonymous):

if you answer my test questions right

OpenStudy (shubhamsrg):

alright,,ans is 3.. hmm,,that costs 6$ !! :D

OpenStudy (shubhamsrg):

@saranya no one gave me any job,,i can be a cook though (paka bahut acha raha hun na) :P

OpenStudy (anonymous):

noo you have to be the first person to answer so you bummed this one

OpenStudy (saranya):

then go n do that :P

OpenStudy (shubhamsrg):

mom threw me out of the kitchen na,,now who'll explain her !! :P

OpenStudy (shubhamsrg):

@gaminoc congrats buddy,,am outa here..enough for today.. ohh wait,,its enough for this post!! :P

OpenStudy (anonymous):

good getchu a facebook if you gonna be just posting haha

OpenStudy (saranya):

:P

OpenStudy (shubhamsrg):

have one already buddy..and am quite quiet there!! (surprisingly enough :P)